In a lottery game, a player picks six numbers from 1 to 27. If the player matches all six numbers, they win 50,000 dollars. Otherwise, they lose $1.

What is the expected value of this game? $
1
Expert's answer
2020-06-30T11:39:46-0400

probability of winning: six out of total 27 numbers:

=

"\frac{6}{27}\frac{5}{26}\frac{4}{25}\frac{3}{24}\frac{2}{23}\frac{1}{22}" = "\frac{1}{296010}"


Probability of losing :

1- "\frac{1}{296010}" ="\frac{296009}{296010}"


We have the following table for x: amount we win or lose and p(x) probability of win or lose


x: 50000 -1

P(x): "\frac{1}{296010}" "\frac{296009}{296010}"


Expected value : "\Sigma x*P(x)"


= "\frac{1}{296010}" *50000 +"\frac{296009}{296010}" *(-1)


= -0.831


answer: expected value of the game = $-0.831